For each pair of functions, find f(g(x)) and g(f(x)).
31)f(x)=3x+5,g(x)=x^2+1
32)f(x)=x^2-5x+2,g(x)=2x
34)f(x)=-2x^2,g(x)=x+4

Answer by stanbon(75887)

For each pair of functions, find f(g(x)) and g(f(x)).
31)f(x)=3x+5,g(x)=x^2+1
f[g(x)} = f[x^2+1] = 3(x^2+1) + 5 = 3x^2 + 3 + 5 = 3x^2 + 8
g[f(x)] = g[3x+5] = (3x+5)^2 + 1 = 9x^2 + 30x + 26
==========================================================
The others follow the same procedure
-------------------------------------
Cheers,
Stan H.